# Evaluating the Ability of Large Language Models

### v1

@@ 1 Introduction
This paper summarises and extends an earlier paper which investigated the ability of Large Language Models (LLMs) to reason about cardinal directions (Cohn and Blackwell, 2024b). That paper introduced two benchmarks: a small dataset of 100 questions designed to test simple factual understanding of cardinal directions, and a more complex one, with 5760 questions, generated from a set of six templates which requires some spatial reasoning to answer correctly. Already in the earlier paper several LLMs were obtaining very good performance on the small dataset, and newer models do even better. Therefore this update concentrates solely on the larger dataset and documents the performance of newer LLMs on it, conducting a more thorough analysis overall.
Many claims (e.g. (Creswell and Shanahan, 2022; Huang and Chang, 2023; Kojima et al., 2022)) have been made since the emergence of Large Language Models (LLMs) as to their ability to reason. Spatial reasoning is of particular interest since not only does it underlie a human’s ability to operate in the physical world, but also because LLMs are not embodied; so the question arises, have they nonetheless acquired an ability to reason about situations which might occur in the real world? This is the question we address here. Spatial reasoning in general encompasses many aspects including topology, distance, and direction. Here, we restrict ourselves to reasoning about cardinal directions (CDs). CDs are important for many reasons, e.g.: (i) successful navigation and route finding/following usually requires a fundamental understanding and ability to reason about CDs: CDs are crucial to success when using a map. Equally, giving directions often relies, at least in part, on the use of CDs. (ii) Understanding the geography of an area depends on understanding the relative CD of one location to another – migration patterns, climate differences and economic variations are often underpinned by CDs. (iii) Weather patterns are often heavily influenced by the direction of the prevailing wind. (iv) CDs often play a critical role in cultural and historical contexts, e.g. the alignment of the pyramids to the four CDs, or in certain languages – e.g. the aboriginal language Guugu Yimithirr has no words for left or right, and spatial information is mainly conveyed using CDs (Haviland, 1998).
@@
In the original paper, Large Reasoning Models (LRMs), in which the models have been trained to perform inference-time reasoning, had not yet emerged, though chain-of-thought prompting (CoT) (Wei et al., 2022) was already popular [^2] We did not explore CoT as we wanted to measure raw LLM performance, independent of any particular prompting strategy; we do not explore CoT prompting in this paper either.. In the intervening months, LRMs have generally become the most performant LLMs, and this paper provides an update on how LRMs perform on our dataset, as well as some other new LLMs, such as GPT-4.5 preview. Since many LLMs now claim to have reasoning capability, we use the strict definition that an LRM is one which reports the use of reasoning tokens in its response to prompts in our benchmark. In this paper we investigate the following LRMs: azure-o1-2024-12-17, azure-o3-mini-2025-01-31, azure-o3-mini-2025-01-31-high, azure-o4-mini-2025-04-16, and deepseek-reasoner. For futher information on LRMs see Xu et al. (2025) who provide a survey covering the emergence and architectures of LRMs.
@@ 2 Related Work
Despite the rapidly growing amount of research into LLMs and their capabilities there has been relatively little devoted specifically to spatial and/or geographic reasoning, and none which has tested their ability to reason about CDs in the way we do here. Of the existing work, we note benchmarks such as StepGame (Li et al., 2024a; Shi et al., 2022) which aims to test an LLM’s ability to correctly determine the spatial relationship between two objects, given the spatial relations between a larger set of objects, and between 1 and 10 reasoning steps are required to correctly determine the result; the direction relationships are not exclusively CDs, but also include, for example “clock face directions” (B is in the three o’clock direction from C). Not surprisingly, performance deteriorates as the required number of steps increases. However, performance increases markedly when an LLM is used to translate from the English specification to a logical representation and symbolic reasoning is used to compute the relationship. It should also be noted that the setting in StepGame is rather artificial compared to the naturalistic settings in our dataset described below. The SpartQA dataset (Mirzaee et al., 2021) is also focused on assessing spatial reasoning, but does not contain any CDs. The bAbI dataset (Weston et al., 2016) has one task which tests CDs understanding, task 19, which contains 1000 training and 1000 test questions: each instance contains 5 facts stating CDs between two objects, and then a question asking about the relation between two of them. Other work (Yamada et al., 2024) has investigated whether LLMs can acquire an understanding of a spatial environment from a turn-by-turn description of a route, with landmarks named at each turn; whilst the LLMs did perform reasonably well, the experiment did not involve any CDs, only left/right and up/down.
There are different kinds of spatial reasoning tasks which can be considered. Relational composition is one of the most studied from a theoretical point of view. A composition table records the results for all combinations of relations in a particular spatial representation such as RCC; investigations (Cohn, 2023; Cohn and Blackwell, 2024a) into various LLMs’ abilities to compute all RCC compositions found reasonable accuracy levels (reduced when relations are anonymised); however RCC is a purely mereotopological calculus with no notion of direction embedded in its semantics.
Some LLMs have been built specifically for geo-applications, but these do not focus on reasoning about directions but rather aspects such as toponym recognition, e.g. Li et al. (2023).
@@ 3 Experimental Design
[…]
Previous work suggests that models with less than about 40B parameters perform poorly at reasoning (Raman et al., 2024). We test state-of-the-art models, favouring larger offerings. However, we did test four smaller models (OLMO 2, DeepSeek R1 7B, LLama 3 7B, and Mistral Small 24B) which were able to run locally on an Apple M2 Pro MacBook using Ollama [^4] https://ollama.com, accessed May 2025. for comparison. All LLM experiments were conducted using the Golem software [^5] https://github.com/RobBlackwell/golem, accessed May 2025.. An advantage of Golem is that all HTTP API requests and responses are logged, providing a comprehensive record of experiments and results in JSONL [^6] https://jsonlines.org, accessed May 2025. format. Unlike some other benchmark studies, all model versions, settings, hyper-parameters, and results are preserved for subsequent analyses, facilitating transparent and reproducible science, as advocated by Burnell et al. (2023).
@@ 3.1 Prompting
We used zero-shot prompting with a system prompt of “You are a helpful assistant. I will give you a question about directions. The answer is either north, south, east, west, north-east, north-west, south-east or south-west. Please only reply with the answer. No yapping.”. We then presented each question in a new chat. We included “No yapping.” since that has been reported as being beneficial in persuading an LLM to be brief in its response (Stifter, 2024).
Where models allow the temperature to be explicitly set (all models except o1, o3-mini, o4-mini and GPT4.5), we set temperature = 0 to try to achieve deterministic answers, although we show (Blackwell et al., 2024) (and in the results below) that this cannot in general be achieved on models run in the cloud.
We use case-insensitive string comparison and remove spurious punctuation and white space before comparing answers. Our prompts ask for cardinal or intercardinal direction answers only, but allow answers such as “The lake is to the west” if the intended answer is “west”, recording such instances where answers do not strictly meet the rubric. We assess performance using accuracy, using prediction intervals (Blackwell et al., 2024) to report variability (where practical).
@@ 4 Results
Figure 1: Accuracy by LLM. Results shaded in blue are from Cohn and Blackwell (2024b). Results with a black border are LRMs. The red dotted line is the guess rate (0.125, since there are eight possible answers). Where possible we use the model names and versions reported in the LLM response, prefixed by the API provider (e.g. openai or azure). Models run locally are prefixed with ollama.
None of the models tested are accurate for all questions (Fig. 1), although all models have higher accuracy than the guess rate. The best performing model in Cohn and Blackwell (2024b) was GPT3.5 Turbo (0.60), but the best performing model now is o1 (0.92), an increase in accuracy of more than 50% in a year. The top 22 results are all commercially-hosted Large Reasoning Models. The best performing locally-hosted model is mistral-small-24b (0.41). In the original paper we only ran models once, and have not conducted further runs here; new models were run more than once to obtain error bars (even with temperature = 0 there can be non determinism (Blackwell et al., 2024)), except when cost precluded this – many LRMs are expensive, e.g. o1 cost \$458 for one run.
We also see slightly different accuracy when accessing the same OpenAI model via the Azure API versus OpenAI API (GPT3.5-Turbo and GPT4-Turbo) and further analysis suggests that these are statistically different (Blackwell et al., 2024).
@@
Cardinal Directions: all models except GPT4.5888Please see our conclusions about GPT4.5 which we suspect may have been trained on our benchmark. had higher overall accuracy with cardinal than intercardinal directions999 It is interesting to note that humans have been reported to be worse at “oblique” directions, rather than the two orthogonal axes (Meng and Qian, 2005; Appelle, 1972) – the so called “oblique effect”. It is suggested to have a neuropsychological basis, though we are not suggesting the effect here is for the same reason. A more likely explanation, in the world of large language models, might be that the intercardinal relation names are composed of two words (in English and many other, but not all languages), and this might create confusion in an LLM. though there were some notably strange results such as with Claude Sonnet where for example, South East had higher accuracy than North (Claude Sonnet 3.5: 0.825 / 0.614, Claude Sonnet 3.7: 0.801 / 0.607). The previously best model, GPT-3.5T still performs very competitively on the cardinal directions but is let down on its performance on the intercardinals, suggesting that the reasoning available in LRMs really helps for intercardinals (though GPT-4.5 which is not a reasoning model has similar performance across both CDs and intercardinals). Deepseek-reasoner and GPT-4.5 have a very similar performance, though with GPT-4.5 slightly outperforming Deepseek-reasoner on the intercardinals, whilst the situation is reversed on the cardinal directions. It is striking how much better Deepseek-reasoner is compared to the base model, Deepseek-chat, from which it was trained. The smallest model, Mistral, which was run on a laptop, is the bottom of the pack, except for north, northeast, and east, yielding an unexplained asymmetry.
@@ 5 Discussion and Conclusions
The development of LLMs is progressing rapidly (though many believe they will never achieve AGI, let alone ever achieve reliable reasoning abilities, at least without a neuro-symbolic component [^10] Li et al. (2024a) shows that the StepGame benchmark can be almost perfectly solved by using an LLM to translate the problem instances into logical form and then using an ASP reasoner to actually solve the problem.).
@@
Although we have not presented further results on newer LLMs on the small dataset in Cohn and Blackwell (2024b) it is worth noting that we did run all the models reported here on the small dataset and most models had an accuracy of 0.9 or greater. One point of interest is that in one case121212“On a hike, a duck pond is to your north and the nearest town is south. What direction is the pond from the town?”, Ollama deepseek-r1:7b (a distilled version of Deepseek’s best reasoning model, R1), we found that with a temperature of 0 it did not terminate (or at least we cut it off after 3 hours on a single prompt); however, when given the same prompt and with the default temperature it terminated very quickly and gave the correct answer.
@@
    Exploration and/or development of prompting strategies (Bhandari, 2024) to improve performance – either using general methods such as chain of thought or tree of thoughts, or spatial-specific ones such as Visualization-of-Thought (Wu et al., 2024).
@@
    Building a comprehensive benchmark for other aspects of spatial reasoning (e.g. topological, distance) and combinations of these; ideally these would be generated programmatically (cf Li et al. (2024b)).
@@
    Consider situations with more than two objects of interest, so that, e.g. compositional reasoning can be tested, and also reasoning about trajectories (cf Weston et al. (2016)).
@@
    Look at CD reasoning in a multi-modal setting, an area that has already received some attention, e.g. Yin et al. (2025).
@@
    Investigate LLM performance in reasoning about the cardinal directions between actual geographic entities – e.g. Fulman et al. (2024) built a 14 question benchmark to investigate whether LLMs display distortions in estimations of intercardinal directions between cities similarly to the errors made by humans (e.g. estimating whether Reno is east or west of San Diego).
@@
The data for all experiments (including the data from Cohn and Blackwell (2024b)) are available in a GitHub repository at http://bit.ly/4e069XO.

### v2

@@ 1 Introduction
This paper summarises and extends an earlier paper which investigated the ability of Large Language Models (LLMs) to reason about cardinal directions (cohn_et_al:LIPIcs.COSIT.2024.28). That paper introduced two benchmarks: a small dataset of 100 questions designed to test simple factual understanding of cardinal directions, and a more complex one, with 5760 questions, generated from a set of six templates which requires some spatial reasoning to answer correctly. Already in the earlier paper several LLMs were obtaining very good performance on the small dataset, and newer models do even better. Therefore this update concentrates solely on the larger dataset and documents the performance of newer LLMs on it, conducting a more thorough analysis overall.
Many claims (e.g. (creswell2022faithful; huang2023reasoning; kojima2022large)) have been made since the emergence of Large Language Models (LLMs) as to their ability to reason. Spatial reasoning is of particular interest since not only does it underlie a human’s ability to operate in the physical world, but also because LLMs are not embodied; so the question arises, have they nonetheless acquired an ability to reason about situations which might occur in the real world? This is the question we address here. Spatial reasoning in general encompasses many aspects including topology, distance, and direction. Here, we restrict ourselves to reasoning about cardinal directions (CDs). CDs are important for many reasons, e.g.: (i) successful navigation and route finding/following usually requires a fundamental understanding and ability to reason about CDs: CDs are crucial to success when using a map. Equally, giving directions often relies, at least in part, on the use of CDs. (ii) Understanding the geography of an area depends on understanding the relative CD of one location to another – migration patterns, climate differences and economic variations are often underpinned by CDs. (iii) Weather patterns are often heavily influenced by the direction of the prevailing wind. (iv) CDs often play a critical role in cultural and historical contexts, e.g. the alignment of the pyramids to the four CDs, or in certain languages – e.g. the aboriginal language Guugu Yimithirr has no words for left or right, and spatial information is mainly conveyed using CDs (haviland1998guugu).
@@
In the original paper, Large Reasoning Models (LRMs), in which the models have been trained to perform inference-time reasoning, had not yet emerged, though chain-of-thought prompting (CoT) (wei2022chain) was already popular [^2] We did not explore CoT as we wanted to measure raw LLM performance, independent of any particular prompting strategy; we do not explore CoT prompting in this paper either.. In the intervening months, LRMs have generally become the most performant LLMs, and this paper provides an update on how LRMs perform on our dataset, as well as some other new LLMs, such as GPT-4.5 preview. Since many LLMs now claim to have reasoning capability, we use the strict definition that an LRM is one which reports the use of reasoning tokens in its response to prompts in our benchmark. In this paper we investigate the following LRMs: azure-o1-2024-12-17, azure-o3-mini-2025-01-31, azure-o3-mini-2025-01-31-high, azure-o4-mini-2025-04-16, and deepseek-reasoner. For futher information on LRMs see xu2025largereasoningmodelssurvey who provide a survey covering the emergence and architectures of LRMs.
@@ 2 Related Work
Despite the rapidly growing amount of research into LLMs and their capabilities there has been relatively little devoted specifically to spatial and/or geographic reasoning, and none which has tested their ability to reason about CDs in the way we do here. Of the existing work, we note benchmarks such as StepGame (li2024advancing; shi2022stepgame) which aims to test an LLM’s ability to correctly determine the spatial relationship between two objects, given the spatial relations between a larger set of objects, and between 1 and 10 reasoning steps are required to correctly determine the result; the direction relationships are not exclusively CDs, but also include, for example “clock face directions” (B is in the three o’clock direction from C). Not surprisingly, performance deteriorates as the required number of steps increases. However, performance increases markedly when an LLM is used to translate from the English specification to a logical representation and symbolic reasoning is used to compute the relationship. It should also be noted that the setting in StepGame is rather artificial compared to the naturalistic settings in our dataset described below. The SpartQA dataset (mirzaee-etal-2021-spartqa) is also focused on assessing spatial reasoning, but does not contain any CDs. The bAbI dataset (weston2016towards) has one task which tests CDs understanding, task 19, which contains 1000 training and 1000 test questions: each instance contains 5 facts stating CDs between two objects, and then a question asking about the relation between two of them. Other work (yamada2024evaluating) has investigated whether LLMs can acquire an understanding of a spatial environment from a turn-by-turn description of a route, with landmarks named at each turn; whilst the LLMs did perform reasonably well, the experiment did not involve any CDs, only left/right and up/down.
There are different kinds of spatial reasoning tasks which can be considered. Relational composition is one of the most studied from a theoretical point of view. A composition table records the results for all combinations of relations in a particular spatial representation such as RCC; investigations (cohn2023evaluation; cohn2024largelanguagemodelsreason) into various LLMs’ abilities to compute all RCC compositions found reasonable accuracy levels (reduced when relations are anonymised); however RCC is a purely mereotopological calculus with no notion of direction embedded in its semantics.
Some LLMs have been built specifically for geo-applications, but these do not focus on reasoning about directions but rather aspects such as toponym recognition, e.g. Li2023GeoLMEL.
@@ 3 Experimental Design
We then exhaustively generated all forms of these questions for all cardinal and intercardinal directions and ten different locomotion types {cycling, driving, hiking, jogging, perambulating, racing, riding, running, unicycling, walking}. Note that in each template, once one of the directions (between “[ ]”) is fixed, then there are only two possibilities for the second direction [^3] Note that for T4, there could potentially be twice the number of choices if one were to consider, e.g. “[west-east]” as well as “[east-west]”. This could conceivably make a difference to an LLM since in one case the agent is walking in the direction specified, and in the other case in the opposite direction.. Following earlier evidence (raman2024steer) that an LLM’s performance can vary depending on whether a question is phrased as using the first or second person, we also generated questions in the first-person (I am), first-person plural (We are), second-person (You are), third-person singular (He is and She is), and third-person plural (They are) forms. This gives us 6 templates $\times$ 10 locomotion forms $\times$ 6 person forms $\times$ 8 directions $\times$ 2 directions-variations = 5760 questions.
Previous work suggests that models with less than about 40B parameters perform poorly at reasoning (raman2024steer). We test state-of-the-art models, favouring larger offerings. However, we did test four smaller models (OLMO 2, DeepSeek R1 7B, LLama 3 7B, and Mistral Small 24B) which were able to run locally on an Apple M2 Pro MacBook using Ollama [^4] https://ollama.com, accessed May 2025. for comparison. All LLM experiments were conducted using the Golem software [^5] https://github.com/RobBlackwell/golem, accessed May 2025.. An advantage of Golem is that all HTTP API requests and responses are logged, providing a comprehensive record of experiments and results in JSONL [^6] https://jsonlines.org, accessed May 2025. format. Unlike some other benchmark studies, all model versions, settings, hyper-parameters, and results are preserved for subsequent analyses, facilitating transparent and reproducible science, as advocated by burnell2023rethink.
@@ 3.1 Prompting
We used zero-shot prompting with a system prompt of “You are a helpful assistant. I will give you a question about directions. The answer is either north, south, east, west, north-east, north-west, south-east or south-west. Please only reply with the answer. No yapping.”. We then presented each question in a new chat. We included “No yapping.” since that has been reported as being beneficial in persuading an LLM to be brief in its response (stifter2024unit).
Where models allow the temperature to be explicitly set (all models except o1, o3-mini, o4-mini and GPT4.5), we set temperature = 0 to try to achieve deterministic answers, although we show (blackwell2024reproduciblellmevaluationquantifying) (and in the results below) that this cannot in general be achieved on models run in the cloud.
We use case-insensitive string comparison and remove spurious punctuation and white space before comparing answers. Our prompts ask for cardinal or intercardinal direction answers only, but allow answers such as “The lake is to the west” if the intended answer is “west”, recording such instances where answers do not strictly meet the rubric. We assess performance using accuracy, using prediction intervals (blackwell2024reproduciblellmevaluationquantifying) to report variability (where practical).
@@ 4 Results
Figure 1: Accuracy by LLM. Results shaded in blue are from cohn_et_al:LIPIcs.COSIT.2024.28. Results with a black border are LRMs. The red dotted line is the guess rate (0.125, since there are eight possible answers). Where possible we use the model names and versions reported in the LLM response, prefixed by the API provider (e.g. openai or azure). Models run locally are prefixed with ollama.
None of the models tested are accurate for all questions (Fig. 1), although all models have higher accuracy than the guess rate. The best performing model in cohn_et_al:LIPIcs.COSIT.2024.28 was GPT3.5 Turbo (0.60), but the best performing model now is o1 (0.92), an increase in accuracy of more than 50% in a year. The top 22 results are all commercially-hosted Large Reasoning Models. The best performing locally-hosted model is mistral-small-24b (0.41). In the original paper we only ran models once, and have not conducted further runs here; new models were run more than once to obtain error bars (even with temperature = 0 there can be non determinism (blackwell2024reproduciblellmevaluationquantifying)), except when cost precluded this – many LRMs are expensive, e.g. o1 cost \$458 for one run.
We also see slightly different accuracy when accessing the same OpenAI model via the Azure API versus OpenAI API (GPT3.5-Turbo and GPT4-Turbo) and further analysis suggests that these are statistically different (blackwell2024reproduciblellmevaluationquantifying).
@@
Cardinal Directions: all models except GPT4.5888Please see our conclusions about GPT4.5 which we suspect may have been trained on our benchmark. had higher overall accuracy with cardinal than intercardinal directions [^9] It is interesting to note that humans have been reported to be worse at “oblique” directions, rather than the two orthogonal axes (MENG20053402; appelle1972perception) – the so called “oblique effect”. It is suggested to have a neuropsychological basis, though we are not suggesting the effect here is for the same reason. A more likely explanation, in the world of large language models, might be that the intercardinal relation names are composed of two words (in English and many other, but not all languages), and this might create confusion in an LLM. though there were some notably strange results such as with Claude Sonnet where for example, South East had higher accuracy than North (Claude Sonnet 3.5: 0.825 / 0.614, Claude Sonnet 3.7: 0.801 / 0.607). The previously best model, GPT-3.5T still performs very competitively on the cardinal directions but is let down on its performance on the intercardinals, suggesting that the reasoning available in LRMs really helps for intercardinals (though GPT-4.5 which is not a reasoning model has similar performance across both CDs and intercardinals). Deepseek-reasoner and GPT-4.5 have a very similar performance, though with GPT-4.5 slightly outperforming Deepseek-reasoner on the intercardinals, whilst the situation is reversed on the cardinal directions. It is striking how much better Deepseek-reasoner is compared to the base model, Deepseek-chat, from which it was trained. The smallest model, Mistral, which was run on a laptop, is the bottom of the pack, except for north, northeast, and east, yielding an unexplained asymmetry.
@@ 5 Discussion and Conclusions
The development of LLMs is progressing rapidly (though many believe they will never achieve AGI, let alone ever achieve reliable reasoning abilities, at least without a neuro-symbolic component101010li2024advancing shows that the StepGame benchmark can be almost perfectly solved by using an LLM to translate the problem instances into logical form and then using an ASP reasoner to actually solve the problem.).
@@
Although we have not presented further results on newer LLMs on the small dataset in cohn_et_al:LIPIcs.COSIT.2024.28 it is worth noting that we did run all the models reported here on the small dataset and most models had an accuracy of 0.9 or greater. One point of interest is that in one case121212“On a hike, a duck pond is to your north and the nearest town is south. What direction is the pond from the town?”, Ollama deepseek-r1:7b (a distilled version of Deepseek’s best reasoning model, R1), we found that with a temperature of 0 it did not terminate (or at least we cut it off after 3 hours on a single prompt); however, when given the same prompt and with the default temperature it terminated very quickly and gave the correct answer.
@@
    Exploration and/or development of prompting strategies (bhandari2024survey) to improve performance – either using general methods such as chain of thought or tree of thoughts, or spatial-specific ones such as Visualization-of-Thought (wu2024visualization).
@@
    Building a comprehensive benchmark for other aspects of spatial reasoning (e.g. topological, distance) and combinations of these; ideally these would be generated programmatically (cf li2024ijcai).
@@
    Consider situations with more than two objects of interest, so that, e.g. compositional reasoning can be tested, and also reasoning about trajectories (cf weston2016towards).
@@
    Look at CD reasoning in a multi-modal setting, an area that has already received some attention, e.g. 10888462.
@@
    Investigate LLM performance in reasoning about the cardinal directions between actual geographic entities – e.g. Fulman01112024 built a 14 question benchmark to investigate whether LLMs display distortions in estimations of intercardinal directions between cities similarly to the errors made by humans (e.g. estimating whether Reno is east or west of San Diego).
@@
The data for all experiments (including the data from cohn_et_al:LIPIcs.COSIT.2024.28) are available in a GitHub repository at http://bit.ly/4e069XO.